# Zero-Knowledge Proofs in Sublinear Space

### v1

@@ 4.3 The Main Equivalence Theorem
    $\mathcal{T}_{C}$ has height $h=O(T/b_{\text{blk}})$ and maximum degree $\Delta\leq\min\{k,\,1{+}r_{\text{reg}}\}=O(k)$. Each $F_{v}$ is computable in $O\!\big{(}b_{\text{blk}}+C_{\text{commit}}(b_{\text{blk}})\big{)}$ time and $O(b_{\text{blk}})$ space, where $C_{\text{commit}}(b)=O(\mathrm{MSM}(b))=O(b)$ with a preprocessed evaluation-basis SRS (conservatively, $O(b\cdot\mathrm{polylog}\,N)$).
@@
Parameters. Height and degree follow from $G_{C}$ (Section 3): $h=B=O(T/b_{\text{blk}})$ and $\Delta\leq\min\{k,\,1{+}r_{\text{reg}}\}$. By Lemma 5, evaluating a block uses $O(b_{\text{blk}})$ space and $O((1{+}r_{\text{reg}})b_{\text{blk}})=O(b_{\text{blk}})$ time. Commitment time is $C_{\text{commit}}(b_{\text{blk}})$ as stated.
@@
Applying Lemma 8 gives $F_{j}(X)=\sum_{i=1}^{B}G_{i,j}(X)$ as polynomials.
@@ 5 The Sublinear-Space Prover
By combining the Tree Evaluation Equivalence (Theorem 4) with the space-efficient Cook–Mertz algorithm (Theorem 3), we construct a sublinear-space zero-knowledge prover. The core idea is to execute the proof generation not by materializing the trace, but by evaluating a implicit cryptographic computation tree.
@@ 5.1 The Streaming Prover Algorithm
We now give a phase-accurate, Fiat–Shamir–compatible schedule that evaluates the implicit tree in sublinear space and preserves the baseline transcript (for linear PCSs). The prover never materializes the full trace, instead making a small number of streaming passes whose memory footprint is $O(\sqrt{T})$ up to a $\log\log$ factor.
@@
We mirror the baseline order of commitments and challenges:
@@
    Phase B (Wire commits, no challenges). For each register $m$ and block $t$, compute $G_{t,m}$ via $\mathsf{EvalBlock}$ (Alg. 2); commit and add into the running $m$-coordinate using $F_{v}$ (Alg. 3). By Lemma 2, the aggregated commitment equals (in distribution) the baseline wire commitment.
@@
    Phase C (Permutation after sampling $(\beta,\gamma)$). Sample $(\beta,\gamma)$ via FS. Re-stream each block to compute the permutation block factor $F_{t}(\beta,\gamma)$ (Lemma 6); multiply into the running accumulator $Z$ and commit to $Z$ blockwise, aggregating linearly to the global $Z$ commitment.
@@
    Phase D (Quotient after sampling $\alpha$). Sample $\alpha$. Compute and commit to $Q$ in coefficient basis by streaming division of $R$ by $Z_{H}$: perform a blocked backward sweep over the coefficients of $R$, emitting coefficient blocks of $Q$ and aggregating their commitments linearly. No materialization of the full coefficient vector is required; the working set is $O(b_{\text{blk}})$.
@@
    Phase E (Openings after sampling $\zeta,\ldots$). Sample evaluation points $(\zeta,\ldots)$. For each polynomial that must be opened at $\zeta$, compute $f(\zeta)$ in a single streaming pass (barycentric in Lagrange basis; Horner in coefficient basis) without storing $f$. Produce the standard batched opening proofs.
Because the commitments in Phases A–D match the baseline (Cor. 1), the Fiat–Shamir challenges have identical distribution, and thus the final transcript distribution is preserved.
@@
3:Phase B (wires): For $t=1$ to $B$ and each register $m\in[k]$: compute $(\text{local},\mathsf{boundary}_{\text{out}},\text{regM\_vals})\leftarrow\mathsf{EvalBlock}(m,t,\cdot)$; form $G_{t,m}$; $\mathsf{com}_{t,m}\leftarrow\mathsf{Commit}(pp,G_{t,m};r_{t,m})$; aggregate into the $m$-coordinate using $F_{v}$
@@
5:Phase C (permutation): For $t=1$ to $B$: compute $F_{t}(\beta,\gamma)$ by re-streaming block $t$; update $Z\leftarrow Z\cdot F_{t}$; commit to $Z$ blockwise and aggregate linearly
@@
7:Phase D (quotient): Stream the coefficients of $R$ and perform a blocked backward division by $Z_{H}$ to emit coefficient blocks of $Q$; commit to each block in coefficient basis and aggregate linearly
@@
The live workspace at any point consists of $O(b_{\text{blk}})$ local state, $O(1)$ accumulators for permutation/lookup factors, and the Cook–Mertz traversal stack of size $O(\log|\mathcal{T}_{C}|\log\log|\mathcal{T}_{C}|)=O(T/b_{\text{blk}})$ up to a $\log\log$ factor (constant arity). Hence total space
$S(T)\ =\ O\!\big{(}b_{\text{blk}}+\tfrac{T}{b_{\text{blk}}}+\lambda\big{)}\ =\ O(\sqrt{T})\ \text{up to a }\log\log\text{ factor}.$
With an evaluation-basis SRS, each block commit costs $O(b_{\text{blk}})$ MSMs; across $kB=\Theta(T/b_{\text{blk}})$ blocks this is $O(T)$ MSMs, plus linear-time streaming for permutation/quotient passes and standard batched openings.
@@ 5.3 Complexity Analysis
$S(T)=O\!\big{(}\sqrt{T}\cdot\log\log T\ +\ \log T\big{)}=O\!\big{(}\sqrt{T}\cdot\log\log T\big{)}\ \text{ space},$
i.e., $O(\sqrt{T})$ up to a $\log\log$ factor.
@@
  • Oracle space. $\mathcal{O}_{\mathrm{func}}$ uses $O(b_{\text{blk}})$ space (Algorithm 3 and Lemma 5); $\mathcal{O}_{\mathrm{children}}$ and $\mathcal{O}_{\mathrm{leaf}}$ add $O(\log T)$ and $O(1)$, respectively. Hence $S_{\mathrm{oracle}}=O(b_{\text{blk}})$.
  • Cook–Mertz space. By Theorem 3 (Section 2),
    $S_{\mathrm{CM}}=O\!\big{(}\Delta\cdot b_{\text{val}}+\log|\mathcal{T}_{C}|\cdot\log\log|\mathcal{T}_{C}|\big{)}.$
    Here $\Delta=O(k)=O(1)$ and $b_{\text{val}}=\Theta(\lambda)$ (a small tuple of commitments and $O(1)$ field elements). The tree size satisfies $\log|\mathcal{T}_{C}|\leq h\cdot\log\Delta=O(h)$ with $h=O(T/b_{\text{blk}})$. Thus
    $S_{\mathrm{CM}}=O\!\big{(}\lambda+(T/b_{\text{blk}})\cdot\log\log|\mathcal{T}_{C}|\big{)}=O\!\big{(}\lambda+T/b_{\text{blk}}\big{)}\ \text{up to a }\log\log\text{ factor}.$
@@
    $S(T)=O\!\big{(}b_{\text{blk}}+\lambda+T/b_{\text{blk}}\big{)}.$
    Choosing $b_{\text{blk}}=\lfloor\sqrt{T}\rfloor$ minimizes the first and third terms, giving
    $S(T)=O\!\big{(}\sqrt{T}+\lambda\big{)}=O\!\big{(}\sqrt{T}+\log T\big{)}=O(\sqrt{T}).$
@@
$\Theta\!\Big{(}\frac{T}{b_{\text{blk}}}\cdot\big{(}b_{\text{blk}}+C_{\text{commit}}(b_{\text{blk}})\big{)}\Big{)}.$
PCS operations contribute only scheme-dependent $\mathrm{poly}(\lambda)$ factors. Substituting $b_{\text{blk}}=\Theta(\sqrt{T})$ yields $O(T\log N)$ with the stated commitments.
@@
Claim. There exists an evaluation schedule such that each node $(m,t)$ is computed exactly once and retained only until its $O(1)$ consumers at layer t+1t{+}1 have consumed its $\mathsf{aux}$; consequently at most a constant number of blocks per layer are live in addition to the Cook–Mertz stack.
Justification. Process the computation graph in increasing $t$; within each layer, any topological order suffices. By first-order locality with read-degree $r_{\text{reg}}=O(1)$, a node $(m,t)$ is needed only by its own successor (m,t+1)(m,t{+}1) and by at most $r_{\text{reg}}$ cross-register consumers at the same layer increment. Once those $O(1)$ consumers have read its $\mathsf{aux}$, $(m,t)$ can be discarded. This yields exactly $kB=\Theta(T/b_{\text{blk}})$ invocations of $\mathcal{O}_{\mathrm{func}}$ without duplication and adds only $O(1)$ blocks of live state beyond the $O(\lambda+T/b_{\text{blk}})$ stack space of Cook–Mertz, preserving the overall $O(\sqrt{T})$ space bound up to a $\log\log$ factor. ∎
@@ 6 Security Preservation
A critical aspect of our construction is that it achieves its space efficiency without compromising security. We prove that the sublinear-space prover inherits completeness, soundness, and zero-knowledge from the underlying PCS-IOP/NIZK protocol (instantiated over the AIR), and that the change of evaluation order (via tree evaluation) leaves transcript distributions unchanged (Theorem 4).
@@ 6.1 Completeness
Completeness follows from correctness at each layer:
@@
    The Tree Evaluation Equivalence (Theorem 4) shows that evaluating the computation tree yields the same $k$ trace-commitments as the standard linear-space prover, together with consistent auxiliary data.
@@
    The Cook–Mertz procedure faithfully evaluates the (implicit) tree to the same root values.
@@
    The final proof $\pi$ is therefore generated from the identical transcript distribution as in the standard prover.
@@ 6.2 Soundness
We give a black-box reduction. $\mathcal{B}$ runs $\mathcal{A}$ while internally simulating the prover’s computations exactly as in $\Pi_{\mathrm{str}}$. By Theorem 4, the distribution of all commitments, openings, and auxiliary values output by $\Pi_{\mathrm{str}}$ is identical to those of $\Pi_{\mathrm{std}}$ on the same inputs and randomness. Therefore any accepting transcript $(x,\pi)$ produced by $\mathcal{A}$ against $\Pi_{\mathrm{str}}$ is an accepting transcript against $\Pi_{\mathrm{std}}$. Hence $\Pr[\mathcal{B}\ \text{wins against}\ \Pi_{\mathrm{std}}]=\Pr[\mathcal{A}\ \text{wins against}\ \Pi_{\mathrm{str}}]=\varepsilon$.
If, moreover, $\Pi_{\mathrm{std}}$ reduces soundness to the PCS binding property (as in standard analyses), then $\mathcal{B}$ composes with that reduction to yield a PCS-binding adversary with essentially the same advantage. No additional loss (such as a factor $\mathrm{poly}(T)$) is introduced by streaming/tree evaluation. ∎

### v2

@@ 4.3 The Main Equivalence Theorem
    $\mathcal{T}_{C}$ has height $h=O(T/b_{\text{blk}})$ and maximum degree $\Delta\leq\min\{k,\,1{+}r_{\text{reg}}\}=O(k)$. Each $F_{v}$ is computable in $O\!\big(k\,b_{\text{blk}}+C_{\text{commit}}(b_{\text{blk}})\big)$ time and $O(b_{\text{blk}})$ space, where $C_{\text{commit}}(b)=O(\mathrm{MSM}(b))=O(b)$ with a preprocessed evaluation-basis SRS (conservatively, $O(b\cdot\mathrm{polylog}\,N)$). With $k=O(1)$ this is $O\!\big(b_{\text{blk}}+C_{\text{commit}}(b_{\text{blk}})\big)$ time.
@@
Parameters. Height and degree follow from $G_{C}$ (Section 3): $h=B=O(T/b_{\text{blk}})$ and $\Delta\leq\min\{k,\,1{+}r_{\text{reg}}\}$. By Lemma 6, evaluating a block uses $O(k\,b_{\text{blk}})$ time and $O(b_{\text{blk}})$ space (collapsing to $O(b_{\text{blk}})$ time for constant $k$). Commitment time is $C_{\text{commit}}(b_{\text{blk}})$ as stated.
@@
Applying Lemma 9 gives $F_{j}(X)=\sum_{i=1}^{B}G_{i,j}(X)$ as polynomial identities.
@@ 5 The Sublinear-Space Prover
By combining the Tree Evaluation Equivalence (Theorem 4) with the space-efficient Cook–Mertz algorithm (Theorem 3), we construct a sublinear-space zero-knowledge prover. The core idea is to execute the proof generation not by materializing the trace, but by evaluating an implicit cryptographic computation tree.
@@ 5.1 The Streaming Prover Algorithm
We now give a phase-accurate, Fiat–Shamir–compatible schedule that evaluates the implicit tree in sublinear space and preserves the baseline transcript (for linear PCSs). The prover never materializes the full trace, instead making a small number of streaming passes whose memory footprint is $O(\sqrt{T})$ up to polylogarithmic factors.
@@
We mirror the baseline order of commitments and challenges. Critical invariant: only the following aggregated per-polynomial commitments are fed into Fiat–Shamir in this exact order: (i) public polynomial commitments, (ii) trace polynomial commitments $\{\mathsf{Commit}(F_{j})\}_{j\in[k]}$, (iii) permutation accumulator commitment (if used), (iv) quotient polynomial commitment $\mathsf{Commit}(Q)$. Block-level commitments are intermediate artifacts and never enter the transcript.
@@
    Phase B (Wire commits, no challenges). For each register $m$ and block $t$, compute $(\text{local},\mathsf{boundary}_{\text{out}},\text{regM\_vals})\leftarrow\mathsf{EvalBlock}$ (Alg. 2); form $G_{t,m}$; produce $\mathsf{Commit}(pp,G_{t,m};r_{t,m})$ and add it into the running $m$-coordinate using $F_{v}$ (Alg. 3). Only the final per-register aggregate commitment (after summing all blocks) is included in the transcript. By Lemma 2, this aggregate equals (in distribution) the baseline wire commitment.
@@
    Phase C (Permutation after sampling $(\beta,\gamma)$). Sample $(\beta,\gamma)$ via FS. Re-stream each block to compute the permutation block factor $F_{t}(\beta,\gamma)$ (Lemma 7); multiply into the running accumulator $Z$. If the baseline protocol commits to the accumulator column $Z$, we commit blockwise and aggregate linearly, emitting only the final aggregate into the transcript; otherwise we maintain $Z$ only as an internal running value for quotient checks.
@@
    Phase D (Quotient after sampling $\alpha$). Sample $\alpha$. See boxed procedure below.
@@
    Phase E (Openings after sampling $\zeta,\ldots$). Sample evaluation points $(\zeta,\ldots)$. For each polynomial and point, compute $f(\zeta)$ in a single streaming pass per polynomial per point (barycentric in Lagrange basis; Horner in coefficient basis) without storing $f$. Produce the standard batched opening proofs, identical in structure to the baseline.
Phase D – Streaming $Q$ in coefficient basis 1. Re-stream trace rows to compute $R(\omega^{i})$ in order (using fixed FS challenges $\alpha,\beta,\gamma$). 2. Feed $R(\omega^{i})$ into a blocked inverse NTT/IFFT to produce coefficient blocks $\{r_{i}\}$ with workspace $O(b_{\text{blk}})$. 3. Run the backward recurrence for division by $X^{N}-c$ on the fly to emit $q$-coefficient blocks. 4. 4a) If baseline commits $Q$ in coefficient basis: Commit each $q$-block and linearly aggregate. 4b) If baseline commits $Q$ in evaluation basis: run a blocked NTT per $q$-block to evaluations, then commit; still aggregate only the per-polynomial commitment into FS.
Because the commitments in Phases A–D that are hashed into FS match the baseline (Cor. 1) and intermediate block-level commitments are not included, the Fiat–Shamir challenges have identical distribution, and thus the final transcript distribution is preserved for linear PCSs.
@@
3:Phase B (wires): For $t=1$ to $B$ and each register $m\in[k]$: compute $(\text{local},\mathsf{boundary}_{\text{out}},\text{regM\_vals})\leftarrow\mathsf{EvalBlock}(m,t,\cdot)$; form $G_{t,m}$; $\mathsf{com}_{t,m}\leftarrow\mathsf{Commit}(pp,G_{t,m};r_{t,m})$; aggregate into the $m$-coordinate using $F_{v}$; do not include block $\mathsf{com}_{t,m}$ in the transcript
@@
5:Phase C (permutation): For $t=1$ to $B$: compute $F_{t}(\beta,\gamma)$ by re-streaming block $t$; update $Z\leftarrow Z\cdot F_{t}$; if committing to $Z$, commit blockwise and aggregate linearly, emitting only the aggregate
@@
7:Phase D (quotient): Execute boxed streaming procedure for $Q$; emit only the aggregate commitment
@@
The live workspace at any point consists of $O(b_{\text{blk}})$ local state, $O(1)$ accumulators for permutation/lookup factors, and the Cook–Mertz traversal stack of size
$O\!\big(\log|\mathcal{T}_{C}|\cdot\log\log|\mathcal{T}_{C}|\big)\;=\;O\!\big(\log(T/b_{\text{blk}})\cdot\log\log(T/b_{\text{blk}})\big)$
for constant arity. Hence total space
$S(T)\ =\ O\!\big(b_{\text{blk}}+\lambda+\log(T/b_{\text{blk}})\log\log(T/b_{\text{blk}})\big).$
With a preprocessed evaluation-basis SRS, each block commit costs $C_{\text{commit}}(b_{\text{blk}})=O(\mathrm{MSM}(b_{\text{blk}}))=O(b_{\text{blk}})$ using Pippenger windowing (SRS tables are not counted as prover workspace); across $kB=\Theta(T/b_{\text{blk}})$ blocks this is $O(T)$ MSMs, plus linear-time streaming for permutation/quotient passes and standard batched openings. The schedule uses a constant number of streaming passes (wires, permutation, quotient, openings), each linear in $T$.
Parallelism caveat.
Parallelizing across $p$ threads must carefully partition work so peak RAM remains $O(\sqrt{T})$; naive parallelization multiplies workspace by $p$.
@@ 5.3 Complexity Analysis
$S(T)=O\!\big(b_{\text{blk}}+\lambda+\log(T/b_{\text{blk}})\log\log(T/b_{\text{blk}})\big)\ \text{ space}.$
Choosing $b_{\text{blk}}=\Theta(\sqrt{T})$ yields
$S(T)=O\!\big(\sqrt{T}+\log T\log\log T\big)=O\!\big(\sqrt{T}\cdot\mathrm{polylog}\,T\big),$
where the $\log T\log\log T$ term is lower-order for practical $T$.
@@
  • Oracle space. $\mathcal{O}_{\mathrm{func}}$ uses $O(b_{\text{blk}})$ space (Algorithm 3 and Lemma 6); $\mathcal{O}_{\mathrm{children}}$ and $\mathcal{O}_{\mathrm{leaf}}$ add $O(\log T)$ and $O(1)$, respectively. Hence $S_{\mathrm{oracle}}=O(b_{\text{blk}}+\log T)$, which is subsumed by the stated bound.
  • Cook–Mertz space. By Theorem 3 and $|\mathcal{T}_{C}|=\Theta(T/b_{\text{blk}})$ at constant arity,
    $S_{\mathrm{CM}}=O\!\big(\Delta\cdot b_{\text{val}}+\log|\mathcal{T}_{C}|\cdot\log\log|\mathcal{T}_{C}|\big)=O\!\big(\lambda+\log(T/b_{\text{blk}})\log\log(T/b_{\text{blk}})\big).$
@@
    $S(T)=O\!\big(b_{\text{blk}}+\lambda+\log(T/b_{\text{blk}})\log\log(T/b_{\text{blk}})\big).$
    Setting $b_{\text{blk}}=\Theta(\sqrt{T})$ gives the claimed bound.
@@
$\Theta\!\Big(\frac{T}{b_{\text{blk}}}\cdot\big(b_{\text{blk}}+C_{\text{commit}}(b_{\text{blk}})\big)\Big),$
multiplied by scheme-dependent $\mathrm{poly}(\lambda)$ factors. Substituting $b_{\text{blk}}=\Theta(\sqrt{T})$ yields the stated bound. The schedule uses a constant number of streaming passes, each linear in $T$, and thus does not increase the asymptotic time.
@@
Claim. There exists a schedule such that each node $(m,t)$ is computed exactly once and retained only until its $O(1)$ consumers at layer t+1t{+}1 have consumed its $\mathsf{aux}$; consequently at most $O(1)$ blocks per layer are live, plus either (i) an $O(1)$ BFS queue or (ii) the Cook–Mertz DFS stack of size $O(\log(T/b_{\text{blk}})\log\log(T/b_{\text{blk}}))$. Justification. Process the computation graph in increasing $t$; within each layer, any topological order suffices. By first-order locality with read-degree $r_{\text{reg}}=O(1)$, a node $(m,t)$ is needed only by its successor (m,t+1)(m,t{+}1) and by at most $r_{\text{reg}}$ cross-register consumers in the next layer. Once those $O(1)$ consumers have read its $\mathsf{aux}$, $(m,t)$ can be discarded. This yields exactly $kB=\Theta(T/b_{\text{blk}})$ invocations of $\mathcal{O}_{\mathrm{func}}$ without duplication and keeps live state within the claimed bounds. ∎
@@ 6 Security Preservation
A critical aspect of our construction is that it achieves its space efficiency without compromising security. We prove that the sublinear-space prover inherits completeness, soundness, and zero-knowledge from the underlying PCS-IOP/NIZK protocol (instantiated over the AIR), and that the change of evaluation order (via tree evaluation) leaves transcript distributions unchanged (Theorem 4) for linear PCSs when the same basis/SRS is used and only aggregate commitments enter Fiat–Shamir. For hash-/FRI-based PCSs, the prover remains sublinear-space but transcripts typically differ; security then follows from the baseline hash-commitment analysis.
@@ 6.1 Completeness
Completeness follows from the chain: block-respecting transformation $\rightarrow$ Tree Evaluation equivalence $\rightarrow$ Cook-Mertz traversal $\rightarrow$ same aggregate commitments $\rightarrow$ same FS challenges $\rightarrow$ baseline completeness. Specifically:
@@
    The Tree Evaluation Equivalence (Theorem 4) shows that evaluating the computation tree yields the same $k$ trace-commitments as the standard linear-space prover, together with consistent auxiliary data; in particular, the multiplicative block factors for permutation and lookup accumulators (Lemmas 7 and 8) ensure streaming reproduces the baseline accumulators.
@@
    The Cook–Mertz procedure faithfully evaluates the (implicit) tree to the same root values with $O(\log|\mathcal{T}_{C}|\log\log|\mathcal{T}_{C}|)$ stack.
@@
    Only aggregate commitments are included in Fiat–Shamir, matching the baseline ordering and encoding; hence the same challenges are derived.
@@ 6.2 Soundness
We give a black-box reduction. The key insight is that any accepting transcript $(x,\pi)$ from $\Pi_{\mathrm{str}}$ is an accepting transcript of $\Pi_{\mathrm{std}}$ under identical FS inputs.
$\mathcal{B}$ runs $\mathcal{A}$ while internally simulating the prover’s computations exactly as in $\Pi_{\mathrm{str}}$. By Theorem 4, the distribution of all aggregate commitments, openings, and auxiliary values output by $\Pi_{\mathrm{str}}$ is identical to those of $\Pi_{\mathrm{std}}$ on the same inputs and randomness. Moreover, because the same aggregate commitments are fed into Fiat–Shamir in the same order and encoding, the derived challenges are identical. Therefore any accepting transcript $(x,\pi)$ produced by $\mathcal{A}$ against $\Pi_{\mathrm{str}}$ is an accepting transcript against $\Pi_{\mathrm{std}}$. Hence $\Pr[\mathcal{B}\ \text{wins against}\ \Pi_{\mathrm{std}}]=\Pr[\mathcal{A}\ \text{wins against}\ \Pi_{\mathrm{str}}]=\varepsilon$.
The reduction preserves advantage and relies on the underlying assumptions of PCS binding and algebraic soundness of randomized checks. No additional loss (such as a factor $\mathrm{poly}(T)$) is introduced by streaming/tree evaluation. ∎